# Contextual Stochastic Optimization for Order Fulfillment Optimization

### v1

@@ 4.4.3 Quantile Regression
Similar to a classification tree, a regression tree partitions the feature space into multiple disjoint regions, but the split criterion is based on the mean squared error (MSE). While regression trees are typically used for generating conditional point predictions, they can be extended to provide distributional forecasts by preserving all observations within each partitioned region, rather than just summary statistics like the mean. Let $\mathcal{R}_{1},\ldots,\mathcal{R}_{R}$ be the regions partitioned by a regression tree. Given covariates $\mathbf{s}$, the region that $\mathbf{s}$ falls into is denoted as $\mathcal{R}(\mathbf{s})$. The conditional distribution of the new sample can be approximated by the empirical CDF:
@@
To refine this empirical CDF, a kernel density estimation method is applied to the observations in $\mathcal{R}(\mathbf{s})$ (Salari et al. 2022).
A similar approach can be extended to random forests, and this idea of QRF was first introduced by Meinshausen and Ridgeway (2006). In QRF, for a new sample with covariates $\mathbf{s}$, the empirical distribution is derived from all regions associated with $\mathbf{s}$ across all trees in the forest. Let $T$ be the total number of regression trees, and let $\mathcal{R}_{1}^{(t)},\ldots,\mathcal{R}_{R}^{(t)}$ be the partitioned regions of the $t$-th tree, where $t\in[T]$. The region that $\mathbf{s}$ falls into in the $t$-th tree is denoted as $\mathcal{R}^{(t)}(\mathbf{s})$. QRF approximates the conditional distribution using the empirical CDF:
@@
To measure the distributional forecasting performance of different learning models, the paper uses two metrics—the Continuous Ranked Probability Score (CRPS) and pinball loss. CRPS evaluates the closeness of the estimated CDF $\hat{F}$ given covariates $\mathbf{s}$ to the observed ground truth value $d$:
@@
Let $\xi$ and $\xi^{\prime}$ be two independent random variables with distribution $\hat{F}(u\mid\mathbf{s})$. The CRPS can alternatively be expressed as:
@@
This equation allows for an approximation of CRPS by sampling from $\hat{F}(u\mid\mathbf{s})$. Pinball loss measures the accuracy of quantile predictions by penalizing errors asymmetrically, depending on whether the predictions are overestimates or underestimates. Let $q$ be a specific quantile. Then, the pinball loss between the observed value $d$ and the predicted value $\hat{d}$ is defined as:
$\text{pinball}(d,\hat{d})=q\max\{d-\hat{d},0\}+(1-q)\max\{\hat{d}-d,0\}.$ (25)
4.4.4 Applying the C-SAA and the C-RO
It is impractical to solve the CSOFP exactly with the learning models just proposed. Indeed, on the one hand, QR models learn continuous distributions, making it computationally intractable to expand the expectation directly. On the other hand, MC-CLF models leverage the fact that the uncertain delivery deviation vector $\tilde{\mathbf{d}}$ has a finite discrete support, which theoretically allows for expansion of the expectation term. However, MC-CLF models face the curse of dimensionality. The number of possible values of $\tilde{\mathbf{d}}$ grow exponentially in $K$ and $L$. This justifies the methodologies proposed in Section 3.
In the following, the C-SAA problems using samples from the conditional distribution estimates provided by the MC-CLF and QR models are referred to as C-SAA-CLF and C-SAA-QR, respectively. Similarly, the C-RO problems are referred to as C-RO-{B,D}-CLF and C-RO-{B,D}-QR, respectively.
@@ 5.1 Data Overview
The industrial partner provides a month-long dataset in 2023, covering more than 1 million online orders for home delivery and over 20,000 unique stock-keeping units (SKUs). Each order is organized by SKUs, resulting in one or more order lines. For an incoming order, there are, on average, 250 to 1,000 possible fulfillment centers across the company’s network to source SKUs from, including mostly physical stores but also regional distribution centers and metro e-commerce centers. Each location is equipped with multiple eligible carriers that are capable of providing different levels of service (LOS), such as Ground, Same-Day, and Overnight shipping. The set of available carriers span from major shipping companies to crowdsourced gig carriers. There are 10 to 30 carrier-LOS pairs appeared in the dataset and throughout the paper, “carrier” refers to one of these specific carrier-LOS pairs.
Each row of the dataset contains transactional details for a single order line, along with the status of the fulfillment network at the time the order was placed. The order information includes the date and time of order placement, the type, dimensions, weight, and quantity of the SKUs, the fulfillment location from which each order line was sourced, the carrier that shipped the order, the customer location, the promised delivery date, and the actual arrival date. The network status data includes the on-hand inventory levels of the requested SKU and the remaining open capacity at various locations. Additionally, the dataset includes carrier-provided shipping costs and planned transit times.
5.1.1 Learning Results
For the evaluation of the contextual learning models, the dataset is split into two parts: orders from the first three weeks (1.3 million order lines) are used for training, while orders from the last week (0.4 million order lines) are reserved for testing and evaluation.
In Table 1, two metrics—log loss and Brier score—are reported to measure the accuracy of probabilistic predictions for the tuned classification models on the test set. Probability calibration is applied to the uncalibrated models that obtained the best metrics. It is notable that both the log loss and Brier scores show improvement after calibration. In addition, Figure 2 illustrates the performance of an example distributional forecast by multi-class classification against its true empirical distribution on the test set. The distribution of the majority class 0 (on-time deliveries) is predicted quite accurately, while the distributions of late and early deliveries are forecasted sufficiently close, albeit with small deviations.
Table 1: Average Log Loss and Brier Score on Test Set
@@ 5.1.1 Learning Results
Figure 2: Comparison of Calibrated Predicted Joint Distribution with True Empirical Distribution for an Example Carrier on Test Set.
Table 2 presents the pinball losses for different quantiles and the CRPS values of the tuned quantile regression models and classification models. The CRPS values are approximated by drawing 1,000 samples to compute Equation (24). It is observed that the MC-CLF models exhibit higher pinball losses compared to the two quantile regression models, likely due to the discrete nature of classification outputs. However, the CRPS of MC-CLF is the lowest among the three models, despite not being explicitly tuned for CRPS.
Additionally, Figure 3 displays the 95% prediction intervals created by the three conditional distribution learning methods for an example carrier. The visualization aligns with the pinball loss results, showing that QRF generally provides the narrowest interval widths. While the intervals generated by MC-CLF cover most of the observed values, their coverage is less effective at the two extremes compared to the two quantile regression methods. The inferior performance of MC-CLF at extreme quantiles is likely due to the discretization of the output space, which limits precision, and the cumulative probability method, which can cause jumps when selecting quantiles. In contrast, regression tree and QRF provide finer-grained predictions, leading to more accurate and narrower intervals at the extremes.
In summary, MC-CLF models excel at providing distributional forecasts, whereas QR models are better at generating prediction intervals for this dataset.
Table 2: Average Pinball Loss and CRPS on Test Set Across All Carriers.
@@
QRF 0.033 0.057 0.263 0.143 0.096 0.402
(a) Regression Tree Prediction Intervals
(b) MC-CLF Prediction Intervals
(c) QRF Prediction Intervals
Figure 3: Sorted Observed Deviation Values versus 95% Prediction Intervals for an Example Carrier.
@@ 5.2 Baseline Models
To compare with the proposed contextual optimization models, three baseline order fulfillment algorithms are introduced below.
@@
    Pure Cost-Driven Greedy Heuristic (Greedy): This greedy heuristic replicates the current practice employed by the industrial partner’s sourcing engine. The algorithm selects the location-carrier pair with the lowest fulfillment cost for each quantity of an SKU. For order $o$, the greedy problem is
    $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\{\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}(c_{\ell}^{fixed}+c_{ik\ell o}^{ship})z_{ik\ell o}\}.$
@@
    Empirical-SAA: This method follows the approach described in Section 3.4 for C-SAA, with the key difference being that the deviation vector samples are drawn from the empirical distribution of deviations in the training set, without incorporating any contextual information. For each sample $n\in[N]$, let $\xi_{k\ell o}^{n}$ denote the $n$-th devaition value sampled from the empirical distribution $P^{E}_{k\ell}$ for carrier-location pair $(k,\ell)$. The corresponding deviation vector for sample $n$ is then expressed as $\bm{\xi}_{o}^{n}=(\xi_{k\ell o}^{n})_{k\in\mathcal{K},\ell\in\mathcal{L}}$. For order $o$, the empirical SAA problem is
    $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\{\frac{1}{N}\sum_{n=1}^{N}g(\mathbf{z}_{o},\bm{\xi}_{o}^{n})\}.$
@@
    Point Predict, Then Optimize (Point-Pred.-Opt.): Unlike the distributional forecasting models mentioned in Section 4.4, typical regression-based ML models aim at estimates only the conditional mean $\mathbb{E}[\tilde{d}_{k\ell}\mid\mathbf{s}_{k\ell o}]$. In this method, conditional mean regression models $h_{k\ell}$ are trained to map covariates to deviation for each $(k,\ell)\in\mathcal{K}\times\mathcal{L}$. This approach transforms CSOFP into a deterministic optimization problem by directly replacing the random parameter $\tilde{\mathbf{d}}_{o}$ with its point prediction $\hat{\mathbf{d}}_{o}=\{h_{k\ell}(\mathbf{s}_{k\ell o})\}_{k\in\mathcal{K},\ell\in\mathcal{L}}$ in the objective function of Problem (18). This method is effective only if the objective function $g(\mathbf{z}_{o},\tilde{\mathbf{d}}_{o})$ is linear in $\tilde{\mathbf{d}}_{o}$, in which case estimating the conditional distribution simplifies to estimating the conditional mean. This is because, by linearity of expectation, $\mathbb{E}_{\tilde{\mathbf{d}}_{o}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s}_{o})}[g(\mathbf{z}_{o},\tilde{\mathbf{d}}_{o})]=g(\mathbf{z}_{o},\mathbb{E}_{\tilde{\mathbf{d}}_{o}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s}_{o})}[\tilde{\mathbf{d}}_{o}])$. However, this linearity does not hold in the proposed objective function due to the asymmetric penalty costs associated with the deviations. So a point prediction may be biased, potentially leading to suboptimal solutions. To facilitate the experiments, several different regression models are trained to minimize MSE, and the model with the smallest MSE in cross-validation is selected for each carrier. See Appendix for the average MSE of different regression models. For order $o$ and point prediction $\hat{\mathbf{d}}_{o}$, the optimization model is
    $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}g(\mathbf{z}_{o},\hat{\mathbf{d}}_{o}).$
5.3 The Experimental Setting
@@
The optimization models in this study were implemented using Gurobi 11.0.0 and Python 3.9. All experiments were conducted on a server equipped with dual-socket Intel Xeon Gold 6226 CPUs, each featuring 24 cores running at 2.7 GHz. The regression and multi-class classification models were developed using the scikit-learn (Pedregosa et al. 2011), XGBoost (Chen and Guestrin 2016), and CatBoost libraries (Prokhorenkova et al. 2018). For quantile regression, the regression tree and quantile random forest models were implemented using the quantile-forest package (Johnson 2024).
For the following experiments, all ML models are trained on orders from the first three weeks, while 5,000 multi-item orders randomly sampled from the last week are used to evaluate the order fulfillment methods. During both training and inference, for multi-SKU orders, features are aggregated by retaining only the features of the SKU with the highest shipping charges and quantities.
@@
One of the difficulties in this application is that the ground truth is not available for options that were not selected. To overcome this limitation, a simulation environment was built and generate the “ground truth” when an algorithm selects an option. The delivery deviations are drawn from a multinomial distribution whose probabilities were estimated by fitting a fine-tuned, calibrated multi-class classification model based on orders in the last week of the dataset. Obviously, this data is not available to the algorithm and only used by the simulator to replicate real-world conditions.
@@
By default, a constant $\beta_{k}=0.5$ discount factor for item batching is considered since the exact values are not provided. Sensitivity analysis are perform on the impact of this factor. The default penalty for early delivery $\gamma^{-}$ is $0.1$ and that for late delivery $\gamma^{+}$ is $10$. The penalty can be understood as the opportunity cost (lost sale cost) for early (late) deliveries, and they are also analyzed with a range of different values. For methods relying on SAA, the default is to have $Q=20$ independent replications each with a sample size of $N_{1}=30$. The obtained solutions are then evaluated with a default sample size of $N_{2}=300$. The default sample size used for C-RO-D is $N=200$. As for C-RO-B, a default of $95\%$ interval is considered. Moreover, a scaling factor $\eta\in[0,1]$ (default of 0.5) determines the uncertainty budget, defined as $B=\eta(KL).$

### v2

@@ 5.3 Tree-Based Quantile Regression
A regression tree partitions the covariate space into disjoint regions to minimize mean squared error (MSE). Although regression trees are often used for point forecasts (e.g., predicting the mean response in each leaf), they can be adapted for distributional forecasting by preserving all observations within each partitioned region. For a new sample with covariates $\mathbf{s}$, let $\mathcal{R}(\mathbf{s})$ be the leaf region into which $\mathbf{s}$ falls. The conditional CDF of the deviation can then be approximated by the empirical distribution of the training samples in that leaf:
@@
A subsequent kernel density estimation applied to $\{d_{m}:\mathbf{s}_{m}\in\mathcal{R}(\mathbf{s})\}$ can further smooth this step‐function CDF (Salari et al. 2022).
QRF extends this approach by pooling the leaf‐level samples across an ensemble of $T$ trees. For each tree $t$, let $\mathcal{R}^{(t)}(\mathbf{s})$ be leaf region containing $\mathbf{s}.$ The QRF estimate of the conditional CDF is
@@
Because this procedure pools ordered observations directly, it naturally preserves the ordinal support and never produces crossing quantiles.
To measure the distributional forecasting performance, the paper uses two metrics—the Continuous Ranked Probability Score (CRPS) and pinball loss. CRPS evaluates the closeness of the estimated CDF $\hat{F}$ given covariates $\mathbf{s}$ to the observed ground truth value $d$
@@
Let $\xi$ and $\xi^{\prime}$ be two independent random variables with distribution $\hat{F}(u\mid\mathbf{s})$. The CRPS can alternatively be expressed as
@@
This equation allows for an approximation of CRPS by sampling from $\hat{F}(u\mid\mathbf{s})$. Pinball loss measures the accuracy of quantile predictions by penalizing errors asymmetrically, depending on whether the predictions are overestimates or underestimates.
@@ 6.1 Data Overview
The industrial partner provides a month-long dataset in 2023, covering more than 1 million online orders for home delivery and over 20,000 unique SKUs. Each order can be sourced from roughly 250 to 1,000 possible fulfillment centers across the company’s network and shipped via one of 10 to 30 distinct carriers.
The target variable in the machine learning problem, delivery deviation, is defined as the difference between the actual delivered date and the promised delivery date. To enhance the quality of the analysis, order lines with extreme deviations were excluded—specifically, those with absolute values greater than 10 days. Such outliers often reflect exceptional circumstances, including returns to sender, lost shipments, or atypical business orders, which can introduce noise and bias into the model.
6.2 Evaluation of Contextual Distributional Learning of Delivery Deviations
This section evaluates the contextual distributional learning models presented in Section 5 for the case study.
6.2.1 Machine Learning Model Settings
For the evaluation of the learning models, the dataset was partitioned into a training set comprising orders from the first three weeks (approximately 1.3 million order lines) and a test set consisting of orders from the final week (approximately 0.4 million order lines). For multi-SKU orders, features were aggregated by retaining only those of the SKU with the highest shipping charge and quantity. The MC-CLF models were implemented using scikit-learn (Pedregosa et al. 2011) and CatBoost (Prokhorenkova et al. 2018), with probability calibration applied to the best‐performing uncalibrated models based on cross‐validated metrics. Quantile regression models, including regression tree and QRF, were implemented via the quantile-forest package (Johnson 2024). All hyperparameters were tuned using three‑fold cross‑validation in Optuna (Akiba et al. 2019): MC‑CLF models were tuned to minimize log loss, whereas quantile regression models were tuned to minimize CRPS.
6.2.2 Learning Performance Analysis
In Table 1, two standard metrics—log loss and Brier score—are reported to measure the accuracy of probabilistic predictions for each tuned MC-CLF model on the test set. Calibration yields a clear improvement: the calibrated classifier achieves the lowest log loss (0.965) and Brier score (0.031), down from 1.050 and 0.034 for the uncalibrated gradient‑boosted tree (GBT‑CLF), and substantially outperforming both the random‑forest classifier (RF‑CLF; 1.593, 0.054) and the simpler baselines (MLR and classification tree).
Figure 2 presents the joint predicted distributions of (a) the calibrated‑CLF, (b) the QRF, and (c) the discretized regression tree, each plotted against the true empirical distribution for orders fulfilled by a representative carrier. The calibrated‑CLF closely reproduces the empirical mass at zero deviation (on‑time deliveries) and tightly follows the tails corresponding to early and late deliveries. The QRF exhibits slightly greater discrepancies, particularly in the majority classes (-1 to 1), while the discretized regression tree shows the largest misalignment from the empirical distribution.
Table 1: Out-of-Sample Average Log Loss and Brier Score Across All Carriers.
@@ 6.2.2 Learning Performance Analysis
(a) Calibrated-CLF
(b) QRF
(c) Regression Tree (discretized)
Figure 2: Out‐of‐Sample Comparison of the Joint Predicted Distribution and True Empirical Distribution for an Example Carrier.
Table 2 reports pinball losses at selected quantiles and CRPS values for the tuned quantile regression and calibrated-CLF models. CRPS was approximated by drawing 1,000 samples to compute Equation (23). The calibrated-CLF model exhibit higher pinball losses than the two quantile regression models, reflecting its inherently discrete output, but nonetheless achieves the lowest CRPS overall, even though it was not directly tuned for this metric. This superior CRPS performance likely results from calibration producing sharper probability estimates, especially for the majority deviation classes.
Additionally, Figure 3 displays the 95% prediction intervals created by the three contextual distribution learning methods for an example carrier. The visualization aligns with the pinball loss results, showing that QRF generally provides the narrowest interval widths. While the intervals generated by calibrated-CLF cover most of the observed values, their coverage is less effective at the two extremes compared to the two QR methods. The inferior performance of calibrated-CLF at extreme quantiles is likely due to the discretization of the output space, which limits precision, and the cumulative probability method, which can cause jumps when selecting quantiles. In contrast, regression tree and QRF provide finer-grained predictions, leading to more accurate and narrower intervals at the extremes.
In summary, MC-CLF models excel at providing distributional forecasts, whereas QR models are better suited to generating prediction intervals on this dataset. Since QRF benefits from ensembling multiple regression trees to reduce variance and improve accuracy and consistently outperform a single regression tree, only QRF is kept for QR in the subsequent CSO evaluations.
Table 2: Out-of-Sample Average Pinball Loss and CRPS Across All Carriers.
@@
(a) Calibrated-CLF Prediction Intervals
(b) QRF Prediction Intervals
(c) Regression Tree Prediction Intervals
Figure 3: Sorted Observed Deviation Values versus 95% Prediction Intervals for an Example Carrier on Out‑of‑Sample Data.
@@ 6.3 Baseline Models
To compare with the proposed contextual optimization models, four baseline order-fulfillment algorithms are introduced below.
@@
    Pure Cost-Driven Greedy Heuristic (Greedy): Mimics the industrial partner’s sourcing engine by selecting, for each unit of an SKU, the carrier–location pair that minimizes the fulfillment cost. Formally, for each order it solves $\min_{\mathbf{z}\in\mathcal{Z}}\{\sum_{i,k,\ell}(c_{\ell}^{fixed}+c_{ik\ell}^{ship})z_{ik\ell}\}.$
@@
    Empirical‑SAA: Follows the SAA framework of Section 4.3, except that each deviation sample is drawn from the overall empirical distribution $P^{E}$ of the training set, without using any contextual information. For each $n\in[N_{1}]$ and every $(k,\ell)\in\mathcal{K}\times\mathcal{L}$, ${}\xi_{k\ell}^{n}\sim P^{E},~{}\bm{\xi}^{n}=\bigl{(}\xi_{k\ell}^{n}\bigr{)}_{k\in\mathcal{K},\,\ell\in\mathcal{L}}.$ The empirical SAA problem is $\min_{\bm{z}\in\mathcal{Z}}\;\frac{1}{N_{1}}\sum_{n=1}^{N_{1}}g\bigl{(}\bm{z},\bm{\xi}^{n}\bigr{)}.$
@@
    C-Empirical-SAA: A slightly smarter SAA baseline that maintains separate empirical distributions $P^{E}_{k}$ for each carrier. For each $n$ and $(k,\ell)$, $\xi_{k\ell}^{n}\sim P^{E}_{k},~{}\bm{\xi}^{n}=\bigl{(}\xi_{k\ell}^{n}\bigr{)}_{k\in\mathcal{K},\,\ell\in\mathcal{L}},$ and it solves the same SAA problem as above.
  4. 4.
    Point Predict–Then–Optimize (PTO): Trains per-carrier regression models $h_{k}(\cdot)$ to estimate the conditional mean deviations for each $(k,\ell)\in\mathcal{K}\times\mathcal{L}$, producing the point‐prediction vector $\hat{\mathbf{d}}=\{h_{k}(\mathbf{s}_{k\ell})\}_{k\in\mathcal{K},\ell\in\mathcal{L}}$. It then solves the deterministic problem $\min_{\bm{z}\in\mathcal{Z}}g\bigl{(}\bm{z},\hat{\mathbf{d}}\bigr{)}.$
    This method is effective only if the objective function $g(\mathbf{z},\tilde{\mathbf{d}})$ is linear in $\tilde{\mathbf{d}}$, in which case estimating the conditional distribution simplifies to estimating the conditional mean. This is because, by linearity of expectation, $\mathbb{E}_{\tilde{\mathbf{d}}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s})}[g(\mathbf{z}_{o},\tilde{\mathbf{d}})]=g(\mathbf{z},\mathbb{E}_{\tilde{\mathbf{d}}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s})}[\tilde{\mathbf{d}}])$. However, $g(\bm{z},\tilde{\mathbf{d}})$ is nonlinear due to asymmetric penalty costs of the deviations, hence replacing the distribution by its mean may bias the solution. For each carrier, the regression model with the lowest cross‐validated mean squared error is selected (see Appendix for details).
6.4 Experimental Setting
This section outlines the computational environment, simulation framework, and parameter configurations used in the experiments.
6.4.1 Implementation Environment
@@
The optimization models in this study were implemented in Python 3.9 using Gurobi 11.0.0, with eight threads allocated to each order instance. A termination criterion of a 1% MIP gap or a 5-minute time limit was applied. All experiments were conducted on a server equipped with dual-socket Intel Xeon Gold 6226 CPUs, each featuring 24 cores running at 2.7 GHz.
6.4.2 Simulation Framework
@@
One of the difficulties in this application is that the ground truth is not available for options that were not selected. To overcome this limitation, a simulation environment was built to generate synthetic “ground truth” when an algorithm selects an option. Specifically, delivery deviations were drawn from a multinomial distribution, with probabilities estimated by fitting a fine-tuned, calibrated multi-class classification model based on orders from the last week of the dataset (testing period). Obviously, this data was not accessible to the algorithm and only used by the simulator to replicate real-world conditions. The training and calibration procedures of the simulator closely followed the methodology outlined in Section 5.2. An evaluation of the simulator is provided in Appendix.
For the following experiments, as mentioned in Section 6.2, all ML models were trained on order data from the first three weeks. Evaluation was conducted on 5,000 multi-item orders, randomly sampled from the final week. To account for variability, solution metrics were computed across 50 random realizations. All reported metrics were normalized by setting the Greedy baseline to 1 for each random seed; true performance values were omitted to due to data confidentiality.
6.4.3 Model Names and Parameters
In the following, the C-SAA problems using samples from the contextual distribution estimates provided by the MC-CLF and QRF models are referred to as C-SAA-CLF and C-SAA-QRF, respectively. Similarly, the C-RO problems are referred to as C-RO-{B,D}-CLF and C-RO-{B,D}-QRF, respectively.
@@ 6.4.3 Model Names and Parameters
By default, a constant $\beta_{k}=0.5$ consolidation discount factor for all $k\in\mathcal{K}$ was considered. The default penalty for early delivery $\gamma^{-}$ was $0.1$ and that for late delivery $\gamma^{+}$ was $10$. For SAA-based methods, the default was to have $Q=20$ independent replications each with a sample size of $N_{1}=30$. The obtained solutions were then evaluated with a default sample size of $N_{2}=300$. The default sample size used for C-RO-D was $N=200$. As for C-RO-B, a default of $95\%$ interval was considered. Moreover, a scaling factor $\eta\in[0,1]$ (default of 0.5) determined the uncertainty budget in C-RO-B, defined as $B=\eta\,|\mathcal{K}|\,|\mathcal{L}|.$
6.5 Overall Results
Table 3 reports the normalized average and worst‐case objective values across 50 random seeds and all order instances. Based on these results, the following observations are drawn.
Average Performance
The C-SAA variants and C-RO-B-CLF method lead to more reduction in the expected objective than other methods. C‑SAA‑CLF achieves the greatest reduction, with a normalized mean of 0.820 $\pm$ 0.003 (–18.0% versus Greedy), followed by C‑SAA‑QRF at 0.844 $\pm$ 0.003 (–15.6%) and C‑RO‑B‑CLF at 0.842 $\pm$ 0.003 (–15.8%). By comparison, Empirical‑SAA (0.962 $\pm$ 0.002) and PTO (0.878 $\pm$ 0.003) offer smaller improvements, and C‑Empirical‑SAA (0.870 $\pm$ 0.003) performs only marginally better than PTO.
Worst-Case Performance
The C-RO-D methods deliver the strongest guarantees under the worst simulated outcome: C‑RO‑D‑CLF attains 0.779 (–22.1%), C‑RO‑D‑QRF 0.787 (–21.3%), and C‑RO‑B‑CLF 0.811 (–18.9%). These three methods uniformly dominate the baseline and other approaches in the worst‐case metric.
Overall, the C-SAA approaches excel in minimizing the expected objective, while the C‑RO‑D variants excel in worst‐case robustness. Notably, C‑RO‑B‑CLF strikes a strong balance between average efficiency and tail‐risk protection.
Table 3: Average and Worst-Case Realized Objective Values Across Different Order Fulfillment Methods.
Method Average Objective (95% CI) Worst-Case Objective
Greedy 1.000 $\pm$ 0.000 1.000
PTO 0.878 $\pm$ 0.003 0.919
Empirical-SAA 0.962 $\pm$ 0.002 1.049
C-Empirical-SAA 0.870 $\pm$ 0.003 0.896
C-SAA-CLF 0.820 $\pm$ 0.003 0.823
C-SAA-QRF 0.844 $\pm$ 0.003 0.842
C-RO-B-CLF 0.842 $\pm$ 0.003 0.811
C-RO-B-QRF 0.861 $\pm$ 0.003 0.824
C-RO-D-CLF 0.928 $\pm$ 0.003 0.779
C-RO-D-QRF 0.914 $\pm$ 0.003 0.787
  • Note: Top 3 methods for each metric are highlighted in bold.